# How much do language models memorize?

### v1

@@ Abstract
We propose a new method for estimating how much a model “knows” about a datapoint and use it to measure the capacity of modern language models. Prior studies of language model memorization have struggled to disentangle memorization from generalization. We formally separate memorization into two components: unintended memorization, the information a model contains about a specific dataset, and generalization, the information a model contains about the true data-generation process. When we completely eliminate generalization, we can compute the total memorization, which provides an estimate of model capacity: our measurements estimate that GPT-style models have a capacity of approximately 3.6 bits per parameter. We train language models on datasets of increasing size and observe that models memorize until their capacity fills, at which point “grokking” begins, and unintended memorization decreases as models begin to generalize. We train hundreds of transformer language models ranging from $500K$ to $1.5B$ parameters and produce a series of scaling laws relating model capacity and data size to membership inference.
@@ 1 Introduction
Studies of language model extraction argue that a data point is memorized if we can induce the model to generate it (Carlini et al., 2023b; Nasr et al., 2023; Schwarzschild et al., 2024). We argue that such generation does not necessarily serve as a proof of memorization. Language models can be coerced to output almost any string (Geiping et al., 2024); hence the fact that a model outputs something is not necessarily a sign of memorization. To address this issue, some researchers have suggested regularizing the input to the language model, such as by limiting its length Schwarzschild et al. (2024) or matching it to the prefix Carlini et al. (2023b) preceding the memorized sentence. However, even with these constraints, memorization cannot be conclusively proven, as the model’s ability to generalize may still be at play. For example, a good language model prompted to add two numbers can output the correct answer without having seen the equation before. In fact, a recent work Liu et al. (2025) shows that some of the instances than were previously thought as memorized do not even exist in the training set and their extractability is a result of generalization. Additionally, verbatim reproduction of a text is not a prerequisite for memorization; a model may still be memorizing specific patterns or sequences, such as every other token, without generating them verbatim.
Given that extraction is neither necessary nor sufficient, defining memorization accurately is a pressing concern. Existing mathematical definitions of memorization, such as those based on membership inference Shokri et al. (2017) and differential privacy Dwork (2006), are defined in the dataset/distribution level. This makes them inadequate for measuring memorization for certain instances (e.g. a particular textbook). Theoretical notions of ’influence’ Feldman (2020); Feldman & Zhang (2020) have been proposed to define memorization at the instance level, but they fall short of meeting our needs. These definitions focus on the capacity of a training algorithm to memorize inputs, whereas our interest lies in understanding how much a specific data point is memorized within a given model. This distinction is crucial, as we are concerned with the memorization properties of a single model, rather than the distribution of models generated by a particular training algorithm.
To bridge this gap, we propose a novel definition of memorization that quantifies the extent to which a model retains information about a specific datapoint. Our approach leverages the concept of compression rate in bits, where a model is considered to have memorized an input if it can be compressed to a significantly shorter encoding in the presence of the model. This framework draws inspiration from Kolmogorov information Kolmogorov (1963) theory and Shannon information Shannon (1948), but can be easily measured in practice using model likelihoods. We tackle the fundamental challenge of distinguishing between memorization and generalization (Prashanth et al., 2024) by decomposing memorization into two distinct components: unintended memorization, which captures the information a model stores about a particular dataset, and generalization, which represents the knowledge a model has acquired about the underlying data-generating process.
@@ 2 Memorization, intended and unintended
Previous works have attempted to define memorization for machine learning models. We aim to provide precise definitions of memorization that meet our criteria, and offer ways to measure it. See Appendix A.3 for a broader discussion on definitions of memorization.
2.1 A statistical view of memorization
@@
Information theory has developed well understood notions of information for random variables. For a random variable $X$, we often use $H(X)$, the entropy of $X$, to define the amount of information present in $X$. Moreover, for two distinct random variables $X,Y$, we can define $X\mid Y$ to be the uncertainty left in $X$ after fixing $Y$. Having defined this quantity, we can now measure mutual information between $X$ and $Y$ by subtracting the leftover information from the total information: $I(X,Y)=H(X)-H(X\mid Y)$.
Now assume we have a machine learning pipeline. We have a prior $\Theta$ on the underlying model that captures our dataset distribution $X$. And we have a learning algorithm $L$ that maps samples from $X$ to a trained model $\hat{\Theta}$. To understand how much information about $X$ is stored in $\hat{\Theta}$, we can use the notion of mutual information:
@@
$\text{mem}_{I}(\hat{\Theta},X,\Theta)=\text{mem}(X,\Theta)-\text{mem}_{U}(X,\hat{\Theta},\Theta)=I(X,\hat{\Theta})-I(X\mid\Theta,\hat{\Theta})$
Now that we have defined our notions of intended and unintended memorization we turn our attention to practically measuring them. Let us first state a proposition that enables measurement of unintended memorization:
@@ 2.2 Measuring unintended memorization with Kolmogorov Complexity
Let $X=(X_{1},\dots,X_{n})$ be an i.i.d, dataset distribution parametrized by ground-truth model $\theta$. Let $L$ be a training algorithm mapping $X$ to $\hat{\Theta}$. Assume $H(\hat{\Theta})=\ell$ and $H(X_{i})=\ell^{\prime}$  [^2] The trained model and each data sample can be presented using $\ell$ and $\ell^{\prime}$ bits respectively. . Then we have $\Big{|}\operatorname*{E}_{\begin{subarray}{c}x\sim X\\ \hat{\theta}\sim L(x)\end{subarray}}\big{[}\text{mem}^{K}_{U}(x_{i},\hat{\theta},\theta)]\big{]}-\text{mem}_{U}(X_{i},\hat{\Theta},\theta)\Big{|}\leq\epsilon.$ for some constant $\epsilon$ independent of $\theta,\ell,\ell^{\prime}$ and $n$. Moreover, we have tail bounds
$\Pr_{\begin{subarray}{c}x\sim X\\ \hat{\theta}\sim L(x)\end{subarray}}\Big{[}\big{|}\Gamma(x,\hat{\theta})\big{|}\geq r+c\Big{]}\leq e^{-2\frac{c^{2}}{\ell\ell^{\prime}}}.$
2.3 Estimating Kolmogorov with likelihoods
Fixing our notion of Kolmogorov memorization, we now describe how we can estimate $H^{K}$ in different setups. Note that exact calculation of Kolmogorov complexity is known to be uncomputable (the decision version of is undecidable). However, we can still approximate it using the best available compression schemes. Below, we summarize how we approximate each term in our definition.
  • $H^{K}(x\mid\hat{\theta})$: Here, $\hat{\theta}$ is the trained target model, which does not necessarily capture the true data distribution. Because compression rate is inherently tied to the likelihood under a predictive model model (Shannon, 1950), we can easily estimate $H^{K}(x\mid\hat{\theta})$ using $p(x\mid\hat{theta})$, the likelihood of $x$ under the target model.
  • $H^{K}(x\mid\hat{\theta},\theta)$: In this case, the compression algorithm has access to both target and reference models. We simply compute $\max\{p(x\mid\hat{\theta},p(x\mid\theta)\}$. In practice, our choice of reference model is a larger model with the same architecture as $\theta$ trained for many steps on a much wider data distribution.
@@ 2.3 Estimating Kolmogorov with likelihoods
Figure 7: Model capacity estimates across different widths and depths in full and half-precision. Doubling precision from bfloat16 to float32 only increases model capacity from $3.51$ to $3.83$ bits-per-parameter.
@@ Results.
We then compute unintended memorization across the entire dataset with respect to the restricted model (Figure LABEL:fig:text-km-total).
@@ 5.2.1 Functional forms
We use a non-linear least squares solver to find optimal values for $c_{1},c_{2},c_{3}$. Solutions found are $c_{1}=1.34$, $c_{2}=-0.034$, and $-33.14$. We plot the scaling laws along with observed data in Figure LABEL:fig:memb-inf-heatmap. Although the sigmoidal function is slightly simplistic (the points do not perfectly fit) our fit produces estimates within $1-2\%$ of observations.
@@ 5.2.2 Validation on larger models
To validate our predictions, we train models with expected membership $F1$ scores of 0.55, 0.75, and 0.95. For model sizes we select GPT-2 small ($125$M params) and GPT-2 XL ($1.5$B params). Using our scaling law, we solve for the dataset size required to get the desired membership inference score for the given model size (see Table 1 for more information). We train models on the estimated dataset size and measure F1 score (Figure LABEL:fig:memb-inf-heatmap). Our predictions are generally within $1.5$ points of the true F1 score; the score is most inaccurate for estimated F1 of 0.75, which is the point where the sigmoid is steepest. In general, the accuracy of our results indicates that our empirical model of membership inference is relatively accurate and provides evidence for why membership inference attacks fail on models trained on extremely large datasets (Das et al., 2024; Duan et al., 2024; Maini et al., 2024).
6 Conclusion
7 Acknowledgements
Appendix A Appendix
A.1 Related Work
Language models and compression.
Language model capacity.
(Arpit et al., 2017) formalize the idea of effective capacity of a model and its training procedure; they also observe that both representation capacity and training time have a strong impact on empirical model capacity. Several other works measure language model capacity in the number of facts or random labels that can be memorized by a network such as an RNN (Collins et al., 2017; Boo et al., 2019) or transformer (Roberts et al., 2020; Heinzerling & Inui, 2021; Allen-Zhu & Li, 2024), sometimes under quantization. A few research efforts (Yun et al., 2019; Curth et al., 2023; Mahdavi et al., 2024; Kajitsuka & Sato, 2024) have developed theoretical estimates for the capacity of different model architectures, although none have yet scaled to multi-layer modern transformers. We are the first to measure a clear upper-bound in model capacity.
Alternative definitions of memorization.
Unintended memorization is deeply related to the many other definitions of memorization proposed in the literature. We provide a detailed comparison in Table LABEL:tab:defns and Section A.3.
A.2 Related Work: Definitions of Memorization
Prior definitions of memorization.
Some of our findings also relate to the discovery of double descent in machine learning (Belkin et al., 2019; Nakkiran et al., 2019) and language modeling (Xia et al., 2023), as well as general discussions of memorization and generalization in deep learning (Zhang et al., 2017; Tänzer et al., 2022).
Here, we discuss other definitions of memorization.
A.3 Other notions of memorization
A.4 Compression with language models beyond arithmetic coding
Shannon (1948) noted that the optimal compression method for a given source is one that assigns codes to symbols such that the average code length approaches the entropy of the source. Arithmetic coding (Pasco, 1977; Rissanen, 1976) is known to be one optimal way to compress text given a distribution over symbols; it was used in (Delétang et al., 2024) to compress text using modern language models.
Although arithmetic coding is known to be optimal for samples generated from the random process of choice, it may still be sub-optimal for cases where the compressed samples are correlated with the choice of random process. Specifically, in language modeling, the training data is highly correlated with the model itself and hence we might need to treat them differently. For instance, we know from previous work that the models behavior on training data points is different from random samples. A large portion of training data can be generated using greedy decoding (Carlini et al., 2023b; Liu et al., 2025) which is a behavior not expected for randomly sampled data. To this end, we design a new compression technique, a generalization of arithmetic coding.
Ensemble compression.
Sampling from language models involve two key parameters $k$ for $top_{k}$ selection and $t$ for temperature. We design a compression method that sets these parameters adaptively. For instance, for cases where we know we can decode the next $100$ tokens in a greedy fashion, we set $k=1$ to reduce the bit length of arithmetic code. Changing the setup of the coding scheme itself requires a new token to be injected and wastes some number of bits, but it could still be beneficial for the code length. Our compression program uses dynamic programming to find the optimal code with injection of these new tokens in the middle of the text. Notably, our algorithm runs in time $O(n*T)$, where $n$ is the number of tokens and $T$ is the number of possible setups (combination of $t$ and $k$) that we allow.
A.5 How reliable are our linear estimates of capacity?
@@
A.6 Additional memorization results
@@
A.7 Comparison of distributions memorized
@@ A.11 Proof of Proposition 4
Now we use this lemma to prove the statement of the Proposition. Let $f$ be a the density function for the joint distribution $(X_{i}\mid\theta,\hat{\Theta})$. That is $f_{i}(x_{i},\hat{\theta})=\Pr[X_{i}=x_{i}\mid\theta\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }\hat{\Theta}=\hat{\theta}]$. Note that this function is independent of $n$ and $\theta$. By definition we have

### v3

@@ Abstract
We propose a new method for estimating how much a model “knows” about a datapoint and use it to measure the capacity of modern language models. We formally separate memorization into two components: unintended memorization, the information a model contains about a specific dataset, and generalization, the information a model contains about the true data-generation process. By eliminating generalization, we can compute the total memorization of a given model, which provides an estimate of model capacity: our measurements estimate that models in the GPT family have an approximate capacity of 3.6 bits-per-parameter. We train language models on datasets of increasing size and observe that models memorize until their capacity fills, at which point “grokking” begins, and unintended memorization decreases as models begin to generalize. We train hundreds of transformer language models ranging from $500K$ to $1.5B$ parameters and produce a series of scaling laws relating model capacity and data size to membership inference.
@@ 1 Introduction
Studies of language model extraction argue that a datapoint is memorized if we can induce the model to generate it (Carlini et al., 2023b; Nasr et al., 2023; Schwarzschild et al., 2024). We argue that such generation does not necessarily serve as a proof of memorization. Language models can be coerced to output almost any string (Geiping et al., 2024); hence the fact that a model outputs something is not necessarily a sign of memorization. To address this issue, some researchers have suggested regularizing the input to the language model, such as by limiting its length (Schwarzschild et al., 2024) or matching it to the prefix (Carlini et al., 2023b) preceding the memorized sentence. However, none of these constraints allow us to distinguish whether a model outputs a string due to memorization or good generalization. For example, a language model prompted to add two numbers can output the answer without having seen the equation before.
To address this issue, we propose a definition of memorization that quantifies the extent to which a model retains information about a specific datapoint. Our approach leverages the concept of compression rate in bits: a model is considered to have memorized an input if the input can be compressed into a shorter encoding when the model is available. This framework draws inspiration from Kolmogorov information theory (Kolmogorov, 1963) and Shannon information theory (Shannon, 1948), but remains practical by estimating information content using model likelihoods. We address the fundamental challenge of distinguishing memorization from generalization (Prashanth et al., 2024) by decomposing memorization into two distinct components: unintended memorization, which captures the information the model retains about a specific dataset, and generalization, which represents the knowledge the model acquires about the underlying data-generating process. This separation is similar to the approach in (Brown et al., 2021), which defines memorization using conditional mutual information between the dataset and the trained model, conditioned on the true concept. However, our notion differs in that it enables this separation at the instance level using algorithmic definitions of information.
@@ 2 Memorization, intended and unintended
Previous works have attempted to define memorization for machine learning models. We aim to provide precise definitions of memorization that meet our criteria, and offer ways to measure it. See Appendix B for a broader discussion on definitions of memorization.
2.1 Warm-up: A statistical view of memorization
@@
We rely on information theory, which has developed well-understood notions of information for random variables. For a random variable $X$, we use $H(X)$, the entropy of $X$, to define the amount of information present in $X$. For two distinct random variables $X,Y$, we can define $X\mid Y$ as the uncertainty left in $X$ after fixing $Y$. With these definitions, we can now measure mutual information between $X$ and $Y$ by subtracting the leftover information from the total information: $I(X,Y)=H(X)-H(X\mid Y)$.
Now suppose we have a machine learning pipeline. We have a prior $\Theta$ on the underlying model that captures our dataset distribution $X$, and a learning algorithm $L$ that maps samples from $X$ to a trained model $\hat{\Theta}$. To understand how much information about $X$ is stored in $\hat{\Theta}$, we can use the notion of mutual information:
@@
$\text{mem}_{I}(X,\hat{\Theta},\Theta)=\text{mem}(X,\hat{\Theta})-\text{mem}_{U}(X,\hat{\Theta},\Theta)=I(X,\hat{\Theta})-I([X\mid\Theta],\hat{\Theta})$
having defined our notions of intended and unintended memorization we turn our attention to practically measuring them. Let us first state a proposition that enables measurement of unintended memorization:
@@
We note that this statistical definition of memorization was first introduced by Brown et al. (2021), where they theoretically looked at the role of unintended memorization in enabling successful learning for certain tasks. However, this preliminary definition is unsuitable for us as we aim to practically measure instance-level memorization. In particular, since we observe only a single trained model and a single input sample, we are unable to define probabilities or conditional probabilities over samples conditioned on the model.
@@ 2.2 Measuring unintended memorization with Kolmogorov Complexity
Let $X=(X_{1},\dots,X_{n})$ be an i.i.d, dataset distribution parametrized by ground-truth model $\theta$. Let $L$ be a training algorithm mapping $X$ to $\hat{\Theta}$. Assume $H(\hat{\Theta})=\ell$ and $H(X_{i})=\ell^{\prime}$  [^2] The trained model and each data sample can be presented using $\ell$ and $\ell^{\prime}$ bits respectively. . Then we have $\Big{|}\operatorname*{E}_{\begin{subarray}{c}x\sim X\\ \hat{\theta}\sim L(x)\end{subarray}}\big{[}\text{mem}^{K}_{U}(x_{i},\hat{\theta},\theta)]\big{]}-\text{mem}_{U}(X_{i},\hat{\Theta},\theta)\Big{|}\leq\epsilon.$ for some constant $\epsilon$ independent of $\ell,\ell^{\prime}$ and $n$.
2.3 Estimating Kolmogorov with compression
Fixing our notion of Kolmogorov memorization, we now describe how we can estimate $H^{K}$ in different setups. Note that exact calculation of Kolmogorov complexity is known to be uncomputable (Kolmogorov, 1965). However, we can still approximate it using the best available compression schemes. These compression schemes could be arbitrary algorithms, e.g. using prompt optimization as in Schwarzschild et al. (2024) or using text prefixes as in Carlini et al. (2023b).
We adopt arithmetic coding as the most natural compression algorithm for language. Arithmetic coding is not only effective for text compression Delétang et al. (2024), but it also allows code lengths to be computed efficiently using model likelihoods. A promising direction for future research is to design compression algorithms specifically tailored to minimize the code length of training data in machine learning models and use it to obtain more accurate estimation of Kolmogorov complexity and memorization. Below, we summarize how we approximate each term in our memorization definition using model likelihoods.
  • $H^{K}(x\mid\hat{\theta})$: Here, $\hat{\theta}$ is the trained target model, which does not necessarily capture the true data distribution. We do not really calculate the compressed code, instead we use fact that the compression rate of arithmetic coding tied to the likelihood of the model (Shannon, 1950). So, we can estimate $H^{K}(x\mid\hat{\theta})$ by the negative log likelihood of $x$ under the target model, $-log(p(x\mid\hat{\theta}))$.
  • $H^{K}(x\mid\hat{\theta},\theta)$: In this case, the compression algorithm has access to both target and reference models. We simply compute $-log(\max\{p(x\mid\hat{\theta}),p(x\mid\theta)\})$. In practice, our choice of reference model is a larger model with the same architecture as $\theta$ trained for many steps on a much wider data distribution.
A curious reader may notice that we began with a likelihood-based notion of memorization, then shifted to a definition grounded in Kolmogorov complexity, and ultimately returned to likelihood to estimate that complexity. We emphasize, however, that the likelihood used in our approximation of Kolmogorov memorization is distinct from the initial likelihood notion. In particular, this likelihood is dependent on the parameters of the decoding algorithm, such as temperature or top-k sampling. More broadly, we note that any compression algorithm can be used to approximate Kolmogorov complexity, and our choice of arithmetic coding is just one instantiation of this broader framework.
Choice of reference model.
In this work we use two reference models to compute $p(x\mid\hat{\theta})$. In our experiments on synthetic random strings (Section 3) we know the exact underlying data distribution and use that as a reference model. In our experiments on text (Section 4) we select $\theta$ to be a large model of the same family, trained on a much larger superset of the training data for $\hat{\theta}$.
@@ 2.3 Estimating Kolmogorov with compression
Table 1: Model capacity estimates across different widths and depths in full and half-precision. Doubling precision from bfloat16 to float32 only increases model capacity from $3.51$ to $3.83$ bits-per-parameter.
@@ 5.2.1 Functional forms
We use a non-linear least squares solver to find optimal values for $c_{1},c_{2},c_{3}$. Solutions found are $c_{1}=1.34$, $c_{2}=-0.034$, and $-33.14$. We plot the scaling laws along with observed data in Figure 7. Although the sigmoidal function is slightly simplistic (the points do not perfectly fit) our fit produces estimates within $1-2\%$ of observations.
@@ 5.2.2 Validation on larger models
To validate our predictions, we train models with expected membership $F1$ scores of 0.55, 0.75, and 0.95. For model sizes we select GPT-2 small ($125$M params) and GPT-2 XL ($1.5$B params). Using our scaling law, we solve for the dataset size required to get the desired membership inference score for the given model size (see Table 2 for more information). We train models on the estimated dataset size and measure F1 score (shown as circles in Figure 7).
Our predictions are generally within $1.5$ points of the true F1 score; the score is most inaccurate for estimated F1 of 0.75, which is the point where the sigmoid is steepest. In general, the accuracy of our results indicates that our empirical model of membership inference is relatively accurate and provides evidence for why membership inference attacks fail on models trained on extremely large datasets (Das et al., 2024; Duan et al., 2024; Maini et al., 2024).
6 Related Work
Language models and compression.
Language model capacity.
Early research on single-layer perceptrons found that single-layer networks can store up to 2 bits-per-parameter (Cover, 1965; Gardner, 1988; Baldi & Hornik, 1989). (Arpit et al., 2017) formalize the idea of effective capacity of a model and its training procedure; they also observe that both representation capacity and training time have a strong impact on empirical model capacity. Several other works measure language model capacity in the number of facts or random labels that can be memorized by a network such as an RNN (Collins et al., 2017; Boo et al., 2019) or transformer (Roberts et al., 2020; Heinzerling & Inui, 2021; Allen-Zhu & Li, 2024), sometimes under quantization. A few research efforts (Yun et al., 2019; Curth et al., 2023; Mahdavi et al., 2024; Kajitsuka & Sato, 2024) have developed theoretical estimates for the capacity of different model architectures, although none have yet scaled to multi-layer modern transformers. (Shwartz-Ziv et al., 2024) also analyze the ‘capacity’ of neural networks on datasets of varying size. We are the first to measure a clear upper-bound in model capacity using per-sample entropy measurements.
Information regularization in learning theory.
Several works have explored the role of mutual information between the input and output of a learning algorithm (Bassily et al., 2018; Haghifam et al., 2020; Steinke & Zakynthinou, 2020). This concept closely relates to the notion of memorization based on Shannon information, discussed in Section 2.1. Some of our findings also relate to the discovery of double descent in machine learning (Belkin et al., 2019; Nakkiran et al., 2019) and language modeling (Xia et al., 2023), as well as general discussions of memorization and generalization in deep learning (Zhang et al., 2017; Tänzer et al., 2022).
Alternative definitions of memorization.
Unintended memorization is deeply related to the many other definitions of memorization proposed in the literature. We provide a detailed comparison in the following subsections.
Prior definitions of memorization.
7 Conclusion
8 Acknowledgements
Appendix A Appendix
A.1 How reliable are our linear estimates of capacity?
@@
A.2 Additional memorization results
@@
A.3 Comparison of distributions memorized
@@ A.7 Proof of Proposition 4
Now we use this lemma to prove the statement of the Proposition. Let $f$ be a the density function for the joint distribution $(X_{i}\mid\theta,\hat{\Theta})$. That is $f_{i}(x_{i},\hat{\theta})=\Pr[X_{i}=x_{i}\mid\theta\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }\hat{\Theta}=\hat{\theta}]$. Note that this function is independent of $n$. By definition we have